OpenStudy (anonymous):

This year the high school basketball team boasted a winning season by winning one more game than it lost. The team played a total of 23 games. How many games did the team win?

OpenStudy (anonymous):

Lost 11 Won 12

OpenStudy (anonymous):

Just keep replacing numbers and you will find it using trial and error method.

OpenStudy (anonymous):

You can also try dividing 23 by 2, you will get 11.5, the answer is 11 and 12.

OpenStudy (anonymous):

Thankyou(:

OpenStudy (anonymous):

Let x be the no of matches lost So the no of matches won is x+1 now x+x+1(total games)=23 2x+1=23 2x=22 x=11(matches lost) matches won = x+1=12

OpenStudy (anonymous):

Of course, if you're doing linear equations, then shayan's answer is more appropriate.